# Spatially focused magnetic hyperthermia: comparison of MRSh and sLLG equations

### v1

@@ IV.2 The MRSh approach
In this subsection, we summarize our results where the MRSh approach is applied to study the spatial focusing ability given for the parallel combination of AC and DC magnetic fields. These results are in complete agreement with previous studies on spatial focusing recent_focused_hyperthermia ; MRSh_theory_focused ; adv_func_mat ; review_mnp . The applied magnetic field is
@@
where $H$ and $Hb_{0}$ stands for the amplitude of the applied and the static fields, the latter is also called the bias field. Based on the numerical solution of the MRSh equation, the SAR can be calculated and in Fig. 1 (which is directly comparable to Fig. 2. in MRSh_theory_focused ), one finds the variation as a function of the magnitude of the static field, $b_{0}$.
Figure 1: (Color online.) Spatial focusing is shown by plotting SAR values as a function of the DC field which is in parallel combination of the AC magnetic field. The lower panel is for lower frequencies suitable for MPI and on the upper panel one finds higher frequencies needed for magnetic hyperthermia. The AC field strength is chosen to be 30 mT.
@@
The spatial focusing ability can be understood by the dynamic hysteresis loops which are given in Fig. 2 (which is in complete agreement with Fig. 3. in MRSh_theory_focused ).
@@
The area of the loop decreases with increasing DC field amplitude and drops to zero when the DC field strength exceeds that of the AC field (see lower panels of Fig. 2). This explains the drop in the SAR values in Fig. 1 with increasing DC field strength. The largest area is found for vanishing DC amplitude, yielding a peak in the SAR values. However, the decrease is much faster for large frequencies compared to low frequencies, i.e. one can observe a broadening of the peak for MPI frequencies. This is because at large frequencies the dynamic hysteresis loops have a much-reduced area, compared to that at the MPI case for the same bias field.
In addition, increasing the frequency leads to the opening of the hysteresis loop. At very high frequencies, the MNPs cannot follow the applied AC field, causing a delay in their magnetization response and thus in the opening of the dynamic hysteresis loop.
@@ IV.3 The stochastic LLG approach
In Fig. 3, we present the results of our sLLG approach, which can be directly compared to Fig. 1.
Figure 3: Results obtained by the sLLG equation (for field strength 1kA/m and frequencies $5\times 10^{7}$ Hz to the top and $10^{6}$ Hz to the bottom). This can be compared to Fig. 1 given by the MRSh approach. One finds similar bell-shapes for the SAR/ILP values.
Similar bell shapes are found for the SAR/ILP values given by the two different methods. We would like to draw the reader’s attention to the broadening of the peak for smaller frequencies used in MPI; see the lower panel of Fig. 3.
@@ V Spatial localization ability
In this section, we use the conclusion of the previous section that the sLLG approach can be used (with the appropriate choice of parameters) to recover the results from the MRSh equation. Thus, here we use the sLLG equation to demonstrate that better spatial focusing can be achieved for image-guided thermal therapy by MPI by the perpendicular (instead of the parallel) combination of AC and DC fields,
@@
In Fig. 5, we show the dynamic hysteresis loop and spatial focusing obtained for large frequencies.
Figure 5: (Color online.) Dynamic hysteresis loop (for parallel combination (13)) and spatial focusing of SAR values for parallel (13) and perpendicular (14) combinations) obtained by the sLLG method for field strength 1kA/m and frequency $5\times 10^{7}$ Hz. One finds no difference in the spatial focusing ability of parallel and perpendicular orientation of AC and DC fields.
@@
In Fig. 6, we plot the dynamic hysteresis loop and spatial focusing obtained for small frequencies.
@@
The increased spatial focusing ability of perpendicular (compared to the parallel) orientation of AC and DC magnetic fields can be explained as follows.
It was shown in superlocal_jpd that the linear response theory is an adequate approximation for relatively large frequencies and small magnetic fields. Accordingly, linear response theory is not applicable at small frequencies (and at large field strengths).
Similar conclusions can be drawn from dynamic hysteresis loops. The shape and area of the loops depend on the parameters of the applied magnetic field. The higher the frequency, the more the linear response is valid and the shape of the loop becomes ellipse-like; see the upper panel of Fig. 5. For low frequencies, one expects the breakdown of the linear response, and the dynamic hysteresis loop deviates from the ellipse-like, see the upper panel of Fig. 6.
@@
Therefore, in the perpendicular case, the spatial focusing ability is much stronger than in the parallel case; see the lower panel of Fig. 6. If the frequency is high and the field strength is small, one expects a linear response and almost identical spatial focusing ability for the two orientations; see the lower panel of Fig. 5 and the lower panel of Fig. 7, where at a value of $b_{0}=0.5$, energy loss occurs in both cases.
@@ VI Summary
In this work, we compared the MRSh and sLLG equations by employing the concept of magnetic and ordinary viscosity with a particular focus on their ability to achieve spatial localization. We demonstrated that the well-known dynamic hysteresis loops and bell-shaped spatially-focused energy loss results obtained from the MRSh approach can be reproduced by the sLLG equation with appropriate choice of parameters. This allows us to draw conclusions about the spatial focusing ability of parallel and perpendicular combination of AC and DC magnetic fields in the Brownian relaxation regime using the sLLG equation. We found that, at low frequencies and at large field amplitudes, the perpendicular orientation of AC and DC magnetic fields provides significantly better spatial focusing ability, which receives important application in image-guided thermal therapy by MPI.
The essence of MPI lies in the application of a selection magnetic field gradient, which defines a small region of low static field, superimposed with an AC magnetic field (with frequency and amplitude lower than those topically used in magnetic hyperthermia). This produces a signal proportional to particle concentration within the low static field region. Thus, with perpendicular AC and DC fields, one can achieve a further improved spatial control of the SAR for MPI.
Based on our findings, we propose the use of perpendicular AC and DC magnetic fields for image-guided thermal therapy by MPI, which requires low frequencies and high field strength.
@@ Acknowledgement
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” is gratefully acknowledged. This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-375) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund. Project no. TKP2021-NKTA-34 has been implemented with the support provided from the National Research, Development and Innovation Fund of Hungary, financed under the TKP2021-NKTA funding scheme. Supported by the University of Debrecen Program for Scientific Publication.

### v2

@@ IV.2 The MRSh approach
In this subsection, we summarize our results where the MRSh approach is applied to study the spatial focusing ability given for the parallel combination of AC and DC magnetic fields. These results are in complete agreement with previous studies on spatial focusing recent_focused_hyperthermia ; MRSh_theory_focused ; adv_func_mat ; review_mnp . The applied magnetic field,
@@
where $H$ and $Hb_{0}$ stands for the amplitude of the applied and the static fields, the latter is also called the bias field. Based on the numerical solution of the MRSh equation, the SAR can be calculated and in Fig. 1 (which is directly comparable to Fig. 2 in MRSh_theory_focused ), one finds the variation as a function of the magnitude of the static field $b_{0}$.
Figure 1: (Color online.) Spatial focusing is shown by plotting SAR values as a function of the DC field which is in parallel combination of the AC magnetic field. The lower panel is for lower frequencies suitable for MPI and on the upper panel one finds higher frequencies needed for magnetic hyperthermia. The AC field strength is chosen to be 30 mT $\sim$ 24 kA/m.
@@
The spatial focusing ability can be understood by the dynamic hysteresis loops which are given in Fig. 2 (which is also in complete agreement with Fig. 3. in MRSh_theory_focused ).
@@
The area of the loop decreases with increasing DC field amplitude and drops to zero when the DC field strength exceeds the AC field (see lower panels of Fig. 2). This explains the drop in the SAR values in Fig. 1 with increasing DC field strength. The largest area is found for vanishing DC amplitude, yielding a peak in the SAR values. However, the decrease is much faster for large frequencies compared to low frequencies, i.e., one can observe a broadening of the peak for MPI frequencies. This is because at large frequencies the dynamic hysteresis loops have a much-reduced area, compared to that at the MPI case for the same bias field.
In addition, the DC field also modifies the shape of the hysteresis loop. Before the magnetization reaches the saturation, it takes a teardrop shape, and of course shifts the entire loop horizontally. This is because the resultant applied magnetic field is shifted towards positive values by the magnitude of the DC field, thus causing the magnetization to be asymmetric, or in case of a strong bias field (stronger than the amplitude of the AC field) to remain in saturation. Moreover, increasing the frequency leads to the opening of the hysteresis loop. At very high frequencies, the MNPs cannot follow the applied AC field, causing a delay in their magnetization response and thus in the opening of the dynamic hysteresis loop.
@@ IV.3 The stochastic LLG approach
In Fig. 3 we present the results of our sLLG approach, which can be directly compared to Fig. 1.
Figure 3: Results obtained by the sLLG equation (for field strength 1 kA/m $\sim$ 1.3 mT and frequencies $5\times 10^{7}$ Hz to the top and $10^{6}$ Hz to the bottom). This can be compared to Fig. 1 given by the MRSh approach. One finds similar bell-shapes for the SAR/ILP values.
Similar bell shapes are found for the SAR/ILP values given by the two different methods. We would like to draw the reader’s attention to the broadening of the peak for smaller frequencies used in MPI, see the lower panel of Fig. 3.
@@ V Spatial localization ability
In this section we use the conclusion of the previous part, that the sLLG approach can be used (with the appropriate choice of parameters) to recover the results from the MRSh equation. Thus, here we use the sLLG equation to demonstrate that better spatial focusing can be achieved for image-guided thermal therapy by MPI in the perpendicular (instead of the parallel) combination of AC and DC fields,
@@
In Fig. 5 we show the dynamic hysteresis loop and spatial focusing obtained for large frequencies.
Figure 5: (Color online.) Dynamic hysteresis loop (for parallel combination (13)) and spatial focusing of SAR values for parallel (13) and perpendicular (14) combinations) obtained by the sLLG method for field strength 1 kA/m $\sim$ 1.3 mT and frequency $5\times 10^{7}$ Hz. One finds no difference in the spatial focusing ability of parallel and perpendicular orientation of AC and DC fields.
@@
In Fig. 6 we plot the dynamic hysteresis loop and spatial focusing obtained for small frequencies.
@@
The increased spatial focusing ability of perpendicular (compared to the parallel) orientation of AC and DC magnetic fields can be explained as follows. It was shown in Ref. superlocal_jpd that the linear response theory is an adequate approximation for relatively large frequencies and small magnetic fields. Accordingly, linear response theory is not applicable at small frequencies (and at large field strengths).
Similar conclusions can be drawn from dynamic hysteresis loops. The shape and area of the loops depend on the parameters of the applied magnetic field. The higher the frequency, the more the linear response is valid and the shape of the loop becomes ellipse-like, see the upper panel of Fig. 5. For low frequencies, one expects the breakdown of the linear response, and the dynamic hysteresis loop deviates from the ellipse-like, see the upper panel of Fig. 6.
@@
Therefore, in the perpendicular case, the spatial focusing ability is much stronger than in the parallel case, see the lower panel of Fig. 6. If the frequency is high and the field strength is small, one expects a linear response and almost identical spatial focusing ability for the two orientations, see the lower panel of Fig. 5 and the lower panel of Fig. 7, where at a value of $b_{0}=0.5$, energy loss occurs in both cases.
@@ VI Summary
In this work, we compared the MRSh and sLLG equations by employing the concept of magnetic and ordinary viscosity with a particular focus on their ability to achieve spatial localization. We demonstrated that the well-known dynamic hysteresis loops and bell-shaped spatially-focused energy loss results obtained from the MRSh approach can be reproduced by the sLLG equation with appropriate choice of parameters.
Thus, it is likely that the results obtained with sLLG are true regardless of the relaxation mechanism. This confirms our previous finding that, in the case of low-frequency and high-amplitude magnetic fields, perpendicular static and variable magnetic fields provide significantly better spatial focusing than parallel arrangements. Based on our results, we therefore recommend the use of perpendicular combination of alternating and static magnetic fields in MPI-based image-guided magnetic hyperthermia, which requires low frequency and high field strength. Our future goals include supporting spatial localization with the MRSh equation and experimental verification. Further development of the setup used for experimental verification could lead to a device that can be used directly in medicine, the concept of which is shown in the schematic diagram Fig. 8.
Figure 8: (Color online.) Schematic diagram of the concept of equipment directly applicable in medicine.
@@ Acknowledgement
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” is gratefully acknowledged. Supported by the EKÖP-25-2 University Research Scholarship Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. Project no. TKP2021-NKTA-34 has been implemented with the support provided from the National Research, Development and Innovation Fund of Hungary, financed under the TKP2021-NKTA funding scheme. Supported by the University of Debrecen Program for Scientific Publication.